OpenStudy (anonymous):

Adrienne had a party at a local bowling alley. She spent for the lane and per person for shoes. Let c be the total cost of the lane and shoes, and let f be the number of friends at her party. Write an equation to represent the problem. Which is the dependent variable? Explain.

OpenStudy (anonymous):

So the cost c = 75 (for the lane) + 5 f (5 per friend for shoes) c=75+5f I think c is the dependant variable because it depends on what f is. c*9 1/8= 255 1/2 divide both sides by 9 1/8 so then you get c= something because the 9 1/8s cancel each other out. Hint for Brandon - work out how many pages he reads per 1 hour for each time and see if they are the same or not.

OpenStudy (anonymous):

:)